OpenStudy (ksaimouli):

These are known: A. Sin(-x)=-Sinx B. Cos(-x)=cosx C.Cos(x+y)= cosxcosy - sinxsiny D.Sin(x+y)= sinxcosy + cosxsiny 1. Sin^2x + cos^2x= 1 (use C and cos0=1) 2. Sin2x=2sinxcosx 3. Cos2x= cos^2x - sin^2x 4. Cos2x= 2cos^2x - 1 5. Cos2x= 1- 2sin^2x 6. Absolute value( Cos x/2)= Sqrt 1+cosx/2 7. Absolute value (Sin x/2) = Sqrt 1- cosx/2

OpenStudy (ksaimouli):

i need 6 only

OpenStudy (anonymous):

ohh i see ! \[\sqrt{1+\cos(x)/2}=\left| \cos(x/2) \right| like this ?\] like this ?

OpenStudy (ksaimouli):